OpenStudy (anonymous):

Solve by completing the square M^2-9/2m=3/2.

OpenStudy (anonymous):

m^2 - 9/2 m - 3/2 = 0 2m^2 - 9m -3 (2m+1)(m-5)+2 =0 (2m+1)(m-5)= -2 2m+1 = -2 m-5 = -2 m = -3/2 m = 3